# UniMoT: Unified Molecule-Text Language Model with Discrete Token Representation

### v1

@@ Appendix D Details and More Results of Experiments
Molecule generation tasks include caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis.
  • Caption-guided molecule generation involves creating molecular structures from textual descriptions, leveraging NLP and cheminformatics to interpret and translate descriptions into chemical structures.
  • Reagent prediction focuses on identifying suitable reagents for given reactants and desired products, optimizing synthetic routes.
  • Forward reaction prediction forecasts probable products from specific reactants and reagents, using knowledge of chemical reactivity.
  • Retrosynthesis deconstructs target molecules into simpler starting materials.
In molecule generation tasks, evaluating the quality of generated molecules involves several metrics that measure different aspects of similarity and validity.
  • Exact Match checks if the generated molecule is identical to the target molecule, offering a stringent criterion for precise replication but potentially overlooking chemically similar variants.
  • The BLEU score [37], adapted from machine translation, measures the overlap of n-grams (short sequences of atoms or bonds) between generated and target molecules, thus assessing partial similarities.
  • Levenshtein Distance [22] evaluates the minimum number of edits needed to transform the generated molecule into the target, providing insight into structural changes required.
  • RDKit [20], MACCS [10], and Morgan [36] Fingerprint Similarities compare the generated and target molecules based on various molecular fingerprinting methods, which capture different aspects of molecular structure and properties.
  • The Validity [19] metric assesses the proportion of chemically valid molecules generated, ensuring that the output consists of plausible chemical structures.
Together, these metrics offer a comprehensive evaluation framework, balancing exact matches with structural and chemical validity.
Appendix E Limitations
While UniMoT demonstrates considerable advancements in unifying molecule and text modalities for comprehensive understanding and generation tasks, several limitations must be acknowledged. Although UniMoT exhibits strong performance in molecule-to-text and text-to-molecule tasks, it has not been extensively tested on more complex molecule generation tasks such as molecule editing, which require precise modifications to molecular structures. Future work could explore extending UniMoT’s capabilities to handle such sophisticated molecular manipulations.
Due to the scarcity of annotated data in the molecular field, the training of UniMoT is less extensive compared to fields like computer vision. This limitation restricts the model’s ability to fully learn and generalize from diverse molecular structures and properties. In contrast, the visual domain benefits from abundant labeled datasets, allowing for more comprehensive training and better performance. Addressing this data scarcity in the molecular domain is crucial for improving UniMoT’s training effectiveness and overall capabilities.
The current empirical evaluations, though extensive, are primarily conducted on standard datasets and benchmarks; expanding the evaluation to a broader array of datasets and real-world scenarios will provide a more comprehensive understanding of the model’s robustness and generalizability.
Appendix F Broader Impacts

### v2

@@ Appendix D Details and More Results of Experiments
Molecule generation tasks include caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. Caption-guided molecule generation involves creating molecular structures from textual descriptions, leveraging NLP and cheminformatics to interpret and translate descriptions into chemical structures. Reagent prediction focuses on identifying suitable reagents for given reactants and desired products, optimizing synthetic routes. Forward reaction prediction forecasts probable products from specific reactants and reagents, using knowledge of chemical reactivity. Retrosynthesis deconstructs target molecules into simpler starting materials. The results of forward reaction prediction and retrosynthesis deconstructs are presented in Table 13.
In molecule generation tasks, evaluating the quality of generated molecules involves several metrics that measure different aspects of similarity and validity. Exact Match checks if the generated molecule is identical to the target molecule, offering a stringent criterion for precise replication but potentially overlooking chemically similar variants. The BLEU score Papineni et al. [2002], adapted from machine translation, measures the overlap of n-grams (short sequences of atoms or bonds) between generated and target molecules, thus assessing partial similarities. Levenshtein Distance Levenshtein and others [1966] evaluates the minimum number of edits needed to transform the generated molecule into the target, providing insight into structural changes required. RDKit Landrum and others [2006], MACCS Durant et al. [2002], and Morgan Morgan [1965] Fingerprint Similarities compare the generated and target molecules based on various molecular fingerprinting methods, which capture different aspects of molecular structure and properties. The Validity Kusner et al. [2017] metric assesses the proportion of chemically valid molecules generated, ensuring that the output consists of plausible chemical structures. Together, these metrics offer a comprehensive evaluation framework, balancing exact matches with structural and chemical validity.
[…]
Forward Reaction Prediction
Llama 0.000 0.020 42.002 0.001 0.002 0.001 0.039
Vicuna 0.000 0.057 41.690 0.007 0.016 0.006 0.059
Mol-Instructions 0.045 0.654 27.262 0.313 0.509 0.262 1.000
InstructMol 0.536 0.967 10.851 0.776 0.878 0.741 1.000
UniMoT 0.611 0.980 8.297 0.836 0.911 0.807 1.000
Retrosynthesis
Llama 0.000 0.036 46.844 0.018 0.029 0.017 0.010
Vicuna 0.000 0.057 46.877 0.025 0.030 0.021 0.017
Mol-Instructions 0.009 0.705 31.227 0.283 0.487 0.230 1.000
InstructMol 0.407 0.941 13.967 0.753 0.852 0.714 1.000
UniMoT 0.478 0.974 11.634 0.810 0.909 0.771 1.000
Table 13: Performance of molecule generation tasks on the Mol-Instructions Fang et al. [2023] benchmark, including forward reaction prediction and retrosynthesis. Bold indicates the best performance, and underline indicates the second best performance.
Appendix E Additional Ablation Studies
LLM Architecture and Adaptation.
We conducted a comparison of molecule captioning performance across various LLM architectures and adaptation strategies, as illustrated in Table 14. Our experiments show that UniMoT performs well across multiple LLM architectures, including Galactica Taylor et al. [2022] and Mistral Jiang et al. [2023] series, demonstrating its robustness and generalizability. The experiments also indicate that scaling up the LLM to 13B or adopting a full fine-tuning (FFT) strategy yields only marginal improvements in performance compared to using Llama-2-7B with LoRA. While larger models and FFT strategy might offer slight gains in performance, they come at a significant cost in terms of efficiency.
Architecture Adaptation BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
Galactica-125M LoRA 28.7 21.5 34.2 21.1 30.3 31.0
Galactica-1.3B LoRA 30.2 22.8 36.0 22.4 32.2 33.2
Mistral-7B LoRA 32.0 24.2 38.0 24.0 34.1 35.2
Llama-2-7B LoRA 31.3 23.8 37.5 23.7 33.6 34.8
Llama-2-7B FFT 32.0 24.6 38.3 24.3 34.7 35.6
Llama-2-13B LoRA 31.8 24.3 38.0 24.1 34.4 35.3
Table 14: Ablation study on the LLM architecture and adaptation strategy.
Models with Comparable Sizes.
We conducted a comprehensive performance comparison between UniMoT and MolCA Liu et al. [2023b] using models of comparable sizes, as detailed in Table 15. The results show that UniMoT consistently outperforms MolCA across various LLM architectures, including Galactica-125M, Galactica-1.3B, and LLaMA-2-7B. This consistent performance highlights the effectiveness of UniMoT in handling molecule-to-text tasks, further validating the superiority of tokenizer-based architecture over adapter-based architecture. The tokenizer-based architecture can achieve better molecule-text alignment through autoregressive molecule-to-text and text-to-molecule pretraining compared to other architectures.
Model BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
MolCA (Galactica-125M) 25.9 17.5 34.4 16.6 23.9 28.5
MolCA (Galactica-1.3B) 28.6 21.3 36.2 21.4 29.7 32.6
MolCA (Llama-2-7B) 28.2 21.0 33.5 20.9 30.0 30.8
UniMoT (Galactica-125M) 28.7 21.5 34.2 21.1 30.3 31.0
UniMoT (Galactica-1.3B) 30.2 22.8 36.0 22.4 32.2 33.2
UniMoT (Llama-2-7B) 31.3 23.8 37.5 23.7 33.6 34.8
Table 15: Performance of UniMoT and MolCA using comparable model sizes on the molecule captioning task using the PubChem dataset.
Query Size.
We also conducted an ablation study to evaluate the performance of UniMoT with different query sizes, as presented in Table 16. The results indicate that increasing the query size leads to improved performance, with the best performance achieves at a query size of 32. However, this larger query size also demands significantly more training time and memory. Therefore, for a more efficient balance between performance and resource consumption, we opt to use a query size of 16, which still offers strong performance while being more computationally feasible.
Architecture Query Size BLEU-2 BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L METEOR
Llama-2-7B 4 25.1 18.3 30.2 18.5 26.1 27.3
Llama-2-7B 8 29.5 21.3 34.5 21.8 30.9 31.5
Llama-2-7B 16 31.3 23.8 37.5 23.7 33.6 34.8
Llama-2-7B 32 32.2 24.9 38.2 24.4 34.7 35.7
Table 16: Performance of UniMoT with different query sizes on the molecule captioning task using the PubChem dataset.
Appendix F Experimental Results with Additional Baselines
We aim to enhance the molecule comprehension and generation experiments in the main text by including additional baselines such as EdgePred Hu et al. [2019], GraphCL You et al. [2020], Mole-BERT Xia et al. [2022], MoMu Su et al. [2022], GIT-Mol Liu et al. [2024b], and MolCA Liu et al. [2023b]. These baselines are presented in Tables 17, 18, 19, and 20.
[…]
KV-PLM 70.50 72.12 55.03 59.83 89.17 54.63 65.40 78.50
EdgePred 67.30 76.00 64.10 60.40 64.10 74.10 76.30 79.90
AttrMask 67.79 75.00 63.57 58.05 75.44 73.76 75.44 80.28
InfoGraph 64.84 76.24 62.68 59.15 76.51 72.97 70.20 77.64
MolCLR 67.79 75.55 64.58 58.66 84.22 72.76 75.88 71.14
GraphMVP 68.11 77.06 65.11 60.64 84.46 74.38 77.74 80.48
GraphCL 69.70 73.90 62.40 60.50 76.00 69.80 78.50 75.40
Mole-BERT 71.90 76.80 64.30 62.80 78.90 78.60 78.20 80.80
MoMu-S 70.50 75.60 63.40 60.50 79.90 70.50 75.90 76.70
MoMu-K 70.10 75.60 63.00 60.40 77.40 71.10 76.20 77.10
MoleculeSTM 69.98 76.91 65.05 60.96 92.53 73.40 76.93 80.77
GIT-Mol 73.90 75.90 66.80 63.40 88.30 - - 81.08
InstructMol (Vicuna-7B) 70.00 74.67 64.29 57.80 91.48 74.62 68.90 82.30
MolCA (OPT-1.3B) 70.00 77.20 64.50 63.00 89.50 - - 79.80
UniMoT (Llama-2-7B) 71.37 76.43 65.78 59.79 92.89 75.97 78.49 83.69
Table 17: ROC-AUC (%) of molecular property prediction task (classification) on the MoleculeNet Wu et al. [2018] datasets. Bold indicates the best performance and underline indicates the second best performance.
[…]
MolT5-Small (T5-Small) 22.5 15.2 30.4 13.5 20.3 24.0
MolT5-Base (T5-Base) 24.5 16.6 32.2 14.0 21.4 26.1
MolT5-Large (T5-Large) 25.9 17.3 34.1 16.4 23.4 28.0
MoMu-Small (T5-Small) 22.9 16.0 31.0 13.7 20.8 24.4
MoMu-Base (T5-Base) 24.7 16.8 32.5 14.6 22.1 27.2
MoMu-Large (T5-Large) 26.3 18.0 34.8 16.9 24.8 28.7
InstructMol (Vicuna-7B) 18.9 11.7 27.3 11.8 17.8 21.3
MolCA (OPT-125M) 25.9 17.5 34.4 16.6 23.9 28.5
MolCA (OPT-1.3B) 28.6 21.3 36.2 21.4 29.7 32.6
3D-MoLM (Llama-2-7B) 30.3 22.5 36.8 22.3 31.2 33.1
UniMoT (Llama-2-7B) 31.3 23.8 37.5 23.7 33.6 34.8
Table 18: Performance (%) of molecule captioning task on the PubChem Kim et al. [2023] dataset. Bold indicates the best performance and underline indicates the second best performance.
[…]
T5-Small 50.1 41.5 60.2 44.6 54.5 53.2
T5-Base 51.1 42.3 60.7 45.1 55.0 53.9
T5-Large 55.8 46.7 63.0 47.8 56.9 58.6
MolT5-Small (T5-Small) 51.9 43.6 62.0 46.9 56.3 55.1
MolT5-Base (T5-Base) 54.0 45.7 63.4 48.5 57.8 56.9
MolT5-Large (T5-Large) 59.4 50.8 65.4 51.0 59.4 61.4
MoMu-Small (T5-Small) 53.2 44.5 - - 56.4 55.7
MoMu-Base (T5-Base) 54.9 46.2 - - 57.5 57.6
MoMu-Large (T5-Large) 59.9 51.5 - - 59.3 59.7
InstructMol (Vicuna-7B) 47.5 37.1 56.6 39.4 50.2 50.9
MolCA (OPT-125M) 61.6 52.9 67.4 53.3 61.5 63.9
MolCA (OPT-1.3B) 63.9 55.5 69.7 55.8 63.6 66.9
UniMoT (Llama-2-7B) 66.4 58.3 72.2 58.4 66.4 70.3
Table 19: Performance (%) of molecule captioning task on the ChEBI-20 Edwards et al. [2022] dataset. Bold indicates the best performance and underline indicates the second best performance.
[…]
Caption-guided Molecule Generation
Llama 0.000 0.003 59.864 0.005 0.000 0.000 0.003
Vicuna 0.000 0.006 60.356 0.006 0.001 0.000 0.001
Mol-Instructions 0.002 0.345 41.367 0.231 0.412 0.147 1.000
MolT5 0.112 0.546 38.276 0.400 0.538 0.295 0.773
UniMoT 0.237 0.698 27.782 0.543 0.651 0.411 1.000
Reagent Prediction
Llama 0.000 0.003 28.040 0.037 0.001 0.001 0.001
Vicuna 0.000 0.010 27.948 0.038 0.002 0.001 0.007
Mol-Instructions 0.044 0.224 23.167 0.237 0.364 0.213 1.000
InstructMol 0.129 0.610 19.664 0.444 0.539 0.400 1.000
UniMoT 0.167 0.728 14.588 0.549 0.621 0.507 1.000
Forward Reaction Prediction
Llama 0.000 0.020 42.002 0.001 0.002 0.001 0.039
Vicuna 0.000 0.057 41.690 0.007 0.016 0.006 0.059
Mol-Instructions 0.045 0.654 27.262 0.313 0.509 0.262 1.000
InstructMol 0.536 0.967 10.851 0.776 0.878 0.741 1.000
UniMoT 0.611 0.980 8.297 0.836 0.911 0.807 1.000
Retrosynthesis
Llama 0.000 0.036 46.844 0.018 0.029 0.017 0.010
Vicuna 0.000 0.057 46.877 0.025 0.030 0.021 0.017
Mol-Instructions 0.009 0.705 31.227 0.283 0.487 0.230 1.000
InstructMol 0.407 0.941 13.967 0.753 0.852 0.714 1.000
UniMoT 0.478 0.974 11.634 0.810 0.909 0.771 1.000
@@ Appendix F Experimental Results with Additional Baselines
Table 20: Performance of molecule generation tasks on the Mol-Instructions Fang et al. [2023] benchmark, including caption-guided molecule generation, reagent prediction, forward reaction prediction, and retrosynthesis. Bold indicates the best performance, and underline indicates the second best performance.
Appendix G Broader Impacts